# SeeDiff: Off-the-Shelf Seeded Mask Generation from Diffusion Models

### v1

@@ Introduction
Founded upon our observations from analysis, we draw connections between the observed characteristics of attention layers and classical seeded segmentation algorithms (Zhang, Zheng, and Cai 2010). Seeded segmentation first finds seeds (initial starting pixels) and expand segmentation masks to pixels that are similar to seeds. With this realization, we propose a new off-the-shelf framework named SeeDiff that draws inspiration from seeded segmentation (Zhang, Zheng, and Cai 2010) to utilize attention maps of Stable Diffusion to generate masks. In particular, we localize a target class object that corresponds to a text, utilizing the cross-attention maps, which quantify text-image correspondences. The coarse responses of the cross-attention maps provide initial seeds, which are then used to find other parts of the object. Such mask region expansion is performed by using self-attention maps that correspond to seeds to find other missing parts of the object. We perform mask region expansion iteratively, as the feature resolution increases, to exploit semantic information from low-resolution features and fine-grained details from high-resolution features. We further refine a mask, using a background mask that can be obtained by mask region expansion from background seeds extracted from the inverted mask.
@@
To minimize human efforts in collecting annotation, a line of works have tried to borrow the data synthesis capability of generative models to generate synthetic datasets consisting of image-mask pairs. Early methods (Voynov, Morozov, and Babenko 2020; Melas-Kyriazi et al. 2022; Zhang et al. 2021; Li et al. 2022) employ generative adversarial networks (GANs) (Goodfellow et al. 2014), where semantic features from intermediate layers are extracted and passed to need-to-train decoders for mask generation. However, these works either require optimization process and can only perform foreground-background separation (Voynov, Morozov, and Babenko 2020; Melas-Kyriazi et al. 2022) or extra decoders (Zhang et al. 2021; Li et al. 2022) that still need to be trained with few images annotated with pixel-level fine-grained masks. Recent works (Li et al. 2023; Wu et al. 2023b, a; Nguyen et al. 2023) have shifted to diffusion-based (Ho, Jain, and Abbeel 2020) text-to-image generative models (e.g., GLIDE (Nichol et al. 2022), Imagen (Saharia et al. 2022), and Stable Diffusion (Rombach et al. 2022)) for its breakthrough in generation performance and easy control over generation via text. While using diffusion models has led to high-quality image-mask generation, they require either a pre-trained object detector (Li et al. 2023) or a pre-trained segmentation network (Wu et al. 2023b); extra modules that need to be trained (Li et al. 2023; Wu et al. 2023b); or extra learning processes (Nguyen et al. 2023; Ma et al. 2023) for refinement. These processes can lead to poor quality mask generation for classes unknown during such processes, thereby being unable to fully exploit the capabilities of text-to-image generative models trained on a very large-scale dataset. Our work, on the other hand, draws inspiration from seeded segmentation, which we claim to be easily transferable to image-mask generation task with off-the-shelf text-guided diffusion models. As a result, we synthesize high-quality image-mask datasets, without the need for pre-trained segmentation/detection networks, extra trainable modules, or optimization/learning processes.

### v2

@@ Introduction
Founded upon our observations from analysis, we draw connections between the observed characteristics of attention layers and classical seeded segmentation algorithms (Zhang, Zheng, and Cai 2010a). Seeded segmentation first finds seeds (initial starting pixels) and expand segmentation masks to pixels that are similar to seeds. With this realization, we propose a new off-the-shelf framework named SeeDiff that draws inspiration from seeded segmentation (Zhang, Zheng, and Cai 2010a) to utilize attention maps of Stable Diffusion to generate masks. In particular, we localize a target class object that corresponds to a text, utilizing the cross-attention maps, which quantify text-image correspondences. The coarse responses of the cross-attention maps provide initial seeds, which are then used to find other parts of the object. Such mask region expansion is performed by using self-attention maps that correspond to seeds to find other missing parts of the object. We perform mask region expansion iteratively, as the feature resolution increases, to exploit semantic information from low-resolution features and fine-grained details from high-resolution features. We further refine a mask, using a background mask that can be obtained by mask region expansion from background seeds extracted from the inverted mask.
@@
To minimize human efforts in collecting annotation, a line of works have tried to borrow the data synthesis capability of generative models to generate synthetic datasets consisting of image-mask pairs. Early methods (Voynov, Morozov, and Babenko 2020; Melas-Kyriazi et al. 2022; Zhang et al. 2021; Li et al. 2022) employ generative adversarial networks (GANs) (Goodfellow et al. 2014), where semantic features from intermediate layers are extracted and passed to need-to-train decoders for mask generation. However, these works either require optimization process and can only perform foreground-background separation (Voynov, Morozov, and Babenko 2020; Melas-Kyriazi et al. 2022) or extra decoders (Zhang et al. 2021; Li et al. 2022) that still need to be trained with few images annotated with pixel-level fine-grained masks. Recent works (Li et al. 2023a; Wu et al. 2023b, a; Nguyen et al. 2023) have shifted to diffusion-based (Ho, Jain, and Abbeel 2020) text-to-image generative models (e.g., GLIDE (Nichol et al. 2022), Imagen (Saharia et al. 2022), and Stable Diffusion (Rombach et al. 2022)) for its breakthrough in generation performance and easy control over generation via text. While using diffusion models has led to high-quality image-mask generation, they require either a pre-trained object detector (Li et al. 2023a) or a pre-trained segmentation network (Wu et al. 2023b); extra modules that need to be trained (Li et al. 2023a; Wu et al. 2023b); or extra learning processes (Nguyen et al. 2023; Ma et al. 2023) for refinement. These processes can lead to poor quality mask generation for classes unknown during such processes, thereby being unable to fully exploit the capabilities of text-to-image generative models trained on a very large-scale dataset. Our work, on the other hand, draws inspiration from seeded segmentation, which we claim to be easily transferable to image-mask generation task with off-the-shelf text-guided diffusion models. As a result, we synthesize high-quality image-mask datasets, without the need for pre-trained segmentation/detection networks, extra trainable modules, or optimization/learning processes.
@@ Acknowledgments
Supplementary Material
SeeDiff: Off-the-Shelf Seeded Mask Generation from Diffusion Models
In this supplementary material, we provide additional context and elaboration on the field related to our research in Sec. S1. Then, in Sec. S2, we conduct additional experiments for our ablation study. Following the experimental setup of our baseline model, DiffuMask, we performed these additional experiments on various datasets using only basic prompts.
S1 Additional Related Works
S1.1 Seeded Segmentation
Semantic segmentation aims to distinguish the semantic area of the input image by performing pixel-level classification. However, obtaining pixel-level annotations necessary for ensuring model performance is a costly endeavor. In contrast, image-level annotations offer a more economical and efficient alternative, though they necessitate the inference of pixel-level object locations due to the absence of explicit object location information. The spatial continuity of semantic labels within the same object implies that pixels in proximity to a discriminative area typically belong to the same object. This characteristic renders seed-based segmentation an effective approach for region delineation through pixel expansion. Classical methods (Zhang, Zheng, and Cai 2010b; Adams and Bischof 1994; Trémeau and Borel 1997) require manual initialization of the seed, followed by segmentation based on pixel similarity, employing graph theory (Boykov and Jolly 2001; Grady 2006) and energy optimization techniques. However, these approaches are limited by their dependence on user-defined seed placement and similarity criteria based on low-level features (e.g. color, intensity, or texture), which can lead to over-segmentation and inaccuracies. To address these limitations, deep learning-based seeded segmentation methods (Kolesnikov and Lampert 2016a; Huang et al. 2018) have been developed. By combining the seeded segmentation method and deep learning, the seed cues generated by the classification network is used as the initial seed to prevent incorrect seed placement, and the improved level of image features is used to evaluate the similarity between pixels and enable more sophisticated expressions. This allows more accurate segmentation.
S1.2 Semantic Segmentation Using Diffusion Models
Diffusion-based models (Rombach et al. 2022) have garnered significant attention in text-to-image generation and have recently been effectively utilized in discriminative computer vision studies. Diffusion models trained on large-scale datasets like LAION-5B (Schuhmann et al. 2022) learn detailed semantic information through the U-Net (Ronneberger, Fischer, and Brox 2015) architecture during the image synthesis process and understand the correspondence between visual pixels and language These models are applied to various segmentation tasks by leveraging their ability to model complex data distributions and utilize internal representations that are associated with high-level semantic concepts during the generation process.
Diffusion models extend the denoising process to generate segmentation mask distributions (Ji et al. 2023; Tan, Wu, and Pi 2023; Amit et al. 2022) based on image conditions through training. Specifically, these models can predict masks for images using a multi-step denoising process (Gu et al. 2024; Wolleb et al. 2021), and Maskdiff (Le et al. 2024) models the distribution of masks and reconstructs the mask from a noise-to-filter perspective. Additionally, the activation at certain diffusion steps effectively captures semantic information (Baranchuk et al. 2022), enabling strong performance in few-shot semantic segmentation. Many studies also focus on extracting internal representations from pretrained diffusion models, as these representations are well-distinguished and interrelated with high/mid-level semantic concepts that can be described in language. Combining the representational power of diffusion models with the classification capabilities of discriminative models (Xu et al. 2023) or analyzing the influence of visual and linguistic information on pixel-space relationships (Tang et al. 2023). By leveraging the features of diffusion models, you can apply them to semantic segmentation in various ways. However, these models still require additional training and Ground Truth data. In contrast, our approach utilizes a pretrained Stable Diffusion model to perform semantic segmentation using the internal information of the diffusion model, without the need for additional training or ground truth data.
S2 Additional Ablation Studies
S2.1 Zero-Shot Semantic Segmentation
| Train Set | mIoU (%) | ||||
| Methods | Type | Categories | Seen | Unseen | Harmonic |
| Manual Mask Supervision | |||||
| ZS3 | real | 15 | 78.0 | 21.2 | 33.3 |
| CaGNet | real | 15 | 78.6 | 30.3 | 43.7 |
| Joint | real | 15 | 77.7 | 32.5 | 45.9 |
| STRICT | real | 15 | 82.7 | 35.6 | 49.8 |
| SIGN | real | 15 | 83.5 | 41.3 | 55.3 |
| ZegFormer | real | 15 | 86.4 | 63.6 | 73.3 |
| Image and Caption only, pre-trained on CC-12M, CC-3M | |||||
| TCL | real | 15+5 | - | - | 83.3 |
| Pseudo Mask Supervision from Model pre-trained on COCO | |||||
| Li et al. [36] (ResNet101) | synthetic | 15+5 | 62.8 | 50.0 | 55.7 |
| Text(Prompt) Supervision | |||||
| DiffuMask (R50) | synthetic | 15+5 | 60.8 | 50.4 | 55.1 |
| DiffuMask (Swin-B) | synthetic | 15+5 | 71.4 | 65.0 | 68.1 |
| SeeDiff (Ours, R50) | synthetic | 15+5 | 63.4 | 61.5 | 62.4 |
| SeeDiff (Ours, Swin-B) | synthetic | 15+5 | 79.9 | 78.6 | 79.3 |
We provide the experimental results on Table S1 for the VOC dataset, where the training and evaluation were conducted in a zero-shot setting with 15 seen classes and 5 unseen classes. The results include the seen mIoU, unseen mIoU, and harmonic mIoU.
S2.2 Additional Evaluation
We evaluate the performance on other dataset ADE-20, using only the basic prompt in the form of ‘a photo of a ’, without any prompt tuning or utilizing domain knowledge of the datasets. Our goal is to demonstrate that synthetic images and masks can be effectively used under these conditions.
| Train Set | Test set | Car | Person | Motorbike | mIoU () |
| DiffuMask | VOC 2012 val | 74.2 | 71.0 | 63.2 | 69.5 |
| SeeDiff (ours) | VOC 2012 val | 75.9 | 67.5 | 66.3 | 69.9 |
| DiffuMask | Cityscapes val | 84.0 | 70.7 | 23.6 | 59.4 |
| SeeDiff (ours) | Cityscapes val | 86.4 | 67.1 | 28.2 | 60.5 |
| Train Set | Number | Backbone | bus | car | person | mIoU () |
| DiffuMask | 60.0 | R50 | 43.4 | 67.3 | 60.2 | 57.0 |
| SeeDiff (ours) | 60.0 | R50 | 55.8 | 73.6 | 68.2 | 65.8 |
| DiffuMask | 60.0 | Swin-B | 72.8 | 73.4 | 62.6 | 69.6 |
| SeeDiff (ours) | 60.0 | Swin-B | 83.4 | 75.1 | 91.0 | 83.2 |
S2.3 Domain Generalization
We evaluate the performance for domain generalization at Table S2. We fix the text input with a basic prompt ‘a photo of a ’. We evaluate the generated dataset on the validation sets of both Pascal VOC 2012 and Cityscapes, using Mask2Former with a ResNet-50 backbone as the evaluator, following the same settings as our baseline, DiffuMask.
| Method | PAMR | mIoU (%) | |
| DiffuMask | ✗ | 66.3 | |
| ✓ | 68.8 | ||
| Dataset Diffusion | ✗ | 69.7 | |
| ✓ | 72.0 | ||
| SeeDiff (Ours) | ✗ | 74.7 | |
| ✓ | 79.6 |
S2.4 Mask Post-Processing
A pre-binarized soft mask (a.k.a. probability map) obtained from attention responses can be coarse and noisy, leading to an uncertainty over how to set a threshold for binarization. Both of our baselines, DiffuMask (Wu et al. 2023b) and Dataset Diffusion (Nguyen et al. 2023), note that an optimal threshold differs for each class. To handle such uncertainty, they both perform mask refinement post-processing. DiffuMask utilizes semantic affinity learning via AffinityNet (Ahn and Kwak 2018) to find an appropriate threshold for each image. Meanwhile, Dataset Diffusion proposes to self-train a model with uncertainty-aware loss function with masks that have uncertainty regions defined by two thresholds. While these methods may have addressed the uncertainty regions, they require extra learning processes with either a pre-trained segmentation network or two hyperparameters. These kinds of extra learning processes can limit generalization performance to other domains. On the other hand, we aim to perform training-free mask refinement post-processing, such that we do not rely on training procedures with a pre-trained network or an extra hyperparameter (other than a standard threshold for binarization). To this end, we employ a pixel-adaptive mask refinement (PAMR) introduced by (Araslanov and Roth 2020a). Similar to AffinityNet (Ahn and Kwak 2018) utilized in DiffuMask (Wu et al. 2023b), PAMR aims to assign the same class labels to nearby pixels that have similar values. To do so, PAMR iteratively updates each soft mask pixel with the convex combination of the pixel values of its neighbors :
where is a kernel function (i.e., similarity function), which is instantiated as a normalized radial basis function (RBF) in PAMR:
where is the locally computed standard deviation of a pixel value. In this work, we apply PAMR on a soft mask obtained from Stable Diffusion, before the binarization. As a result, we obtain an enhanced mask quality, revealed by the performance gain with PAMR as demonstrated in Tab. S4. One may argue that our final performance is largely due to PAMR. However, Tab. S4 shows that our SeeDiff without PAMR outperforms PAMR-enhanced versions of DiffuMask and Dataset Diffusion.
We also note that our method benefits substantially more from PAMR, compared to other methods. We believe this is because PAMR acts as pixel-level region expansion, similar to our SA-based region expansion. Specifically, PAMR finds similar regions at pixel level, akin to how self-attention finds similar regions, however at patch level. Our SA-based region expansion is restricted to resolution, due to the design of Stable Diffusion, being unable to perform pixel-level region expansion with SA. Thus, a pixel-level region expansion with PAMR as our last step of iterative region expansion can be beneficial and thus considered as a integral component to our framework that takes an iterative procedure of finding similar regions with increasing resolutions.
S2.5 Qualitative Results
Effect of Seed Region Expansion
Fig. S1 presents qualitative results across various images, demonstrating how our method, seeds guided region expansion, compensates for the objectness may missed by the cross-attention aggregation map.
Effect of Mask Refinement with Background Expansion
Fig. S2 presents qualitative results across various images, demonstrating how our method, mask refinement with background expansion, that refine the coarse map, using the high resolution of fine-grained self-attention feature map. Fig. S2 shows the visualization of the final mask before applying the PAMR algorithm, which is part of the mask refinement process. We qualitatively observe that the background expansion effectively refines the coarse mask across images from various classes.
Comparison with Previous Works : Basic Prompt
Fig. S3 shows the comparison with previous works, DiffuMask and Dataset Diffusion. In the case of Diffumask setting, the final map was derived by aggregating the cross-attention maps across all resolutions. For Dataset Diffusion, the final map was generated by first producing the cross-attention aggregation map at a resolution of 16 and the self-attention aggregation map at a resolution of 32, then multiplying these maps and applying exponentiation using the value of . We applied PAMR consistently in the mask post-processing stage across all methods to compare the masks generated under the same conditions using the basic prompt ‘a photo of a ’ without any prompt tuning. Additionally, we uniformly set the threshold parameter for generating the binary mask input to PAMR at .
Open-Vocabulary Mask Generation
Not only our methodology but also previous works utilize Stable Diffusion as the backbone model, given its excellent generalization performance. This strong generalization capability allows the model to generate masks for the target class even when various sentences with different expressions are used as input text in an open-vocabulary setting. Therefore, we qualitatively demonstrate the robustness of mask generation for the target class, even with various text inputs in Fig. S4. And also, in Fig. S5, we visualize our generated mask with various open-vocabulary text prompts.